OpenStudy (anonymous):

Xavier is riding on a Ferris wheel at the local fair. His height can be modeled by the equation H(t) = 20 cos pi over 15 + 30, where H represents the height of the person above the ground in feet at t seconds.

OpenStudy (anonymous):

Assuming Xavier begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Xavier's height above the ground reaches a minimum?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

Is the equation missing the value t?

OpenStudy (anonymous):

Sorry, Its supposed to say (pi/15)t

OpenStudy (anonymous):

LIke this ? \[H (t) = 20 \cos (\frac{ \pi }{ 15 } t)\]

OpenStudy (anonymous):

\[H(t)=20\cos [(\pi/15)*t]+30\]

OpenStudy (anonymous):

Yeah, like that!

OpenStudy (anonymous):

But + 30

OpenStudy (anonymous):

Ok H(t)= 20 cos ((pi/15)t) + 30 You know that the graph of cos pi fluctuates between 1 and -1 in the y direction, with -1 being the lowest point. So guess what t would make cos the lowest? Separate cos (pi/15) (t) = -1 pi/15 t = cos ^-1 (-1) pi/15 t = pi t = 15 Plug that back in the equation and find H(t) H(t) = 10

OpenStudy (anonymous):

the graph of cos (x) *

OpenStudy (anonymous):

THANK YOU!!!!!!!!!